# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ Conclusion.
Let $f(x_{i})=\frac{\mu(x_{i}+\left\|\mathbf{x}_{-i}\right\|)^{\gamma}x_{i}}{x_{i}+(1+\alpha)\left\|\mathbf{x}_{-i}\right\|}-c_{i}(x_{i})$ be the utility function for player $i$ under no sharing. Assume we have a profile $\mathbf{x}$ such that $\left\|\mathbf{x}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$, where $\mathcal{X}(\rho)$ is the ESE in $G(\rho)$.
Let $Y_{LB}=\min_{k\in[n]}\sum_{i\neq k}x_{i}^{\min}$ and $Y_{UB}=\sum_{j=1}^{n}x_{j}^{\max}-\min_{k}x_{k}^{\max}$, with the condition that $0.5Y_{LB}>\delta\sqrt{n}$. Then, $f(x_{i})$ is Lipschitz continuous in $x_{i}$ with Lipschitz constant at most:
$L$ $=\mu\Bigg{[}\frac{\gamma(0.5Y_{LB})^{\gamma-1}(x_{i}^{\max})^{2}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{\gamma(0.5Y_{LB})^{\gamma-1}x_{i}^{\max}(1+\alpha)(1.5Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(x_{i}^{\max}+1.5Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{(x_{i}^{\max}+1.5Y_{UB})^{\gamma}(1+\alpha)(1.5Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(1.5Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}\Bigg{]}+c_{i}^{\prime}(x_{i}^{\max}).$
Proof of Lemma 23.
First, we establish bounds on $\left\|\mathbf{x}_{-i}\right\|$ for the perturbed profile $\mathbf{x}$. While we know that $Y_{LB}\leq\left\|\mathcal{X}(\rho)_{-i}\right\|\leq Y_{UB}$ for the ESE, we need to determine bounds for our perturbed profile.
By the triangle inequality and the relationship between norms:
$\left|\left\|\mathbf{x}_{-i}\right\|-\left\|\mathcal{X}(\rho)_{-i}\right\|\right|$ $\leq\left\|\mathbf{x}_{-i}-\mathcal{X}(\rho)_{-i}\right\|$
$\leq\sqrt{n-1}\left\|\mathbf{x}_{-i}-\mathcal{X}(\rho)_{-i}\right\|_{2}$
$\leq\sqrt{n}\left\|\mathbf{x}-\mathcal{X}(\rho)\right\|_{2}$
$\leq\sqrt{n}\delta.$
Therefore:
$\left\|\mathbf{x}_{-i}\right\|$ $\geq\left\|\mathcal{X}(\rho)_{-i}\right\|-\sqrt{n}\delta\geq Y_{LB}-\sqrt{n}\delta$
$\left\|\mathbf{x}_{-i}\right\|$ $\leq\left\|\mathcal{X}(\rho)_{-i}\right\|+\sqrt{n}\delta\leq Y_{UB}+\sqrt{n}\delta.$
Given our assumption that $0.5Y_{LB}>\delta\sqrt{n}$, we have:
$\left\|\mathbf{x}_{-i}\right\|$ $>Y_{LB}-0.5Y_{LB}=0.5Y_{LB}>0$
@@
$\left\|\mathbf{x}_{-i}\right\|$ $\leq Y_{UB}+\sqrt{n}\delta<Y_{UB}+0.5Y_{LB}\leq 1.5Y_{UB}.$
@@
Let $y=\left\|\mathbf{x}_{-i}\right\|$ for simplicity. We compute the derivative of $f(x_{i})$ using the quotient rule:
@@
Using our established bounds $y\geq 0.5Y_{LB}$, $y\leq 1.5Y_{UB}$, and $x_{i}\leq x_{i}^{\max}$:
$\left|f^{\prime}(x_{i})\right|$ $\leq\mu\Bigg{[}\frac{\gamma(0.5Y_{LB})^{\gamma-1}(x_{i}^{\max})^{2}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{\gamma(0.5Y_{LB})^{\gamma-1}x_{i}^{\max}(1+\alpha)(1.5Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(x_{i}^{\max}+1.5Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{(x_{i}^{\max}+1.5Y_{UB})^{\gamma}(1+\alpha)(1.5Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(1.5Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}\Bigg{]}+c_{i}^{\prime}(x_{i}^{\max}).$
Therefore, $f(x_{i})$ is Lipschitz continuous with constant $L$ as claimed. ∎
C.2 Complete Proof of Theorem 6
See 6
Proof of Theorem 6.
@@ C.2 Complete Proof of Theorem 6
Let us start by laying out all the notations we need for the proof.
  • $X_{UB}=\sum_{i=1}^{n}x_{i}^{\max}$ denotes upper bound on the total quality in ESE, and $Y_{UB}=\sum_{j=1}^{n}x_{j}^{\max}-\min_{k}x_{k}^{\max}$ denotes upper bound on the total quality in ESE excluding some single player (notice that this bound is not player-specific). They hold due to Lemma 10.
  • $X_{LB}=\sum_{i=1}^{n}x_{i}^{\min}$ denotes lower bound on the total quality in ESE, and $Y_{LB}=\min_{k\in[n]}\sum_{i\neq k}x_{i}^{\min}$ denotes lower bound on the total quality in ESE excluding some single player (notice that this bound is not player-specific). They hold due to Lemma 14.
  • $L=\frac{\mu\sqrt{n}}{\beta}\frac{X_{UB}}{(X_{LB})^{2-\gamma}}$ is the Lipschitz constant of $\mathcal{X}(\cdot)$, as shown in Proposition 16.
  • $M_{1}=M_{c^{\prime}}+\mu\cdot B(0.5Y_{LB},1.5Y_{UB},n)$ is the constant from Proposition 20, where $M_{c^{\prime}}$ is the maximum of the derivatives of the cost functions, and $B(Y_{LB},Y_{UB},n)$ is the constant from Lemma 21, defined as:
    $B(Y_{LB},Y_{UB},n)$ $=(Y_{LB})^{\gamma-3}\Big{[}x^{\max}Y_{UB}(1+(4+\alpha)\sqrt{n}+2Y_{LB})$
    $+Y_{UB}^{2}+\sqrt{n}(1+\alpha)(x^{\max})^{2}\Big{]}.$
    where $x^{\max}=\max_{i\in[n]}x_{i}^{\max}$.
  • $M_{2}=\mu\left(\gamma(0.5Y_{LB})^{\gamma-1}\sqrt{n}+(1.5Y_{UB})^{\gamma}\right)$ is the constant from Proposition 22.
@@
Before starting the proof, we address a special case of the choice of a large $\varepsilon$. Although we don’t expect the algorithm to be executed with such a large $\varepsilon$, we want to show that the algorithm will prevail in this case as well. Notice that if $\varepsilon>\mu X_{UB}$, then no matter which output the algorithm returns, it will satisfy the approximation guarantees:
@@
  • First, notice that the utilities of the players are bounded by $\mu X_{UB}$. Thus, no player can have a deviation that is larger than $X_{UB}$. Thus, the algorithm will return an $\varepsilon$-FSE.
@@
  • In the same line of reasoning, the platform revenue is also bounded by $\mu X_{UB}$, thus, the algorithm will return a solution with revenue not smaller than $\text{OPT}-\varepsilon$.
@@
Given these reasons, we limit our attention to the case when $\varepsilon\leq\mu X_{UB}$, as we just showed that the other case is trivial.
Part 1: Approximation guarantees.
@@
First, let us specify the choice of the constant $\delta$ in the algorithm. We set $\delta=\frac{\varepsilon}{A}$, where $A=\max\left\{\frac{\sqrt{n}\mu X_{UB}}{0.5Y_{LB}},4M_{1}(1+L),4M_{2}(1+L)\right\}$. The reasons for this choice will become clear in the next paragraphs.
Let $\rho^{\star}\in\mathcal{F}$ be a feasible solution (induces an FSE) such that $U_{P}(\mathcal{X}(\rho^{\star}),\mathbf{1};f_{\rho^{\star}})\geq\text{OPT}-\frac{1}{2}\varepsilon$. Our objective is to show that Algorithm 1 returns a parameter $\hat{\rho}$ such that $(\mathcal{X}(\hat{\rho}),\mathbf{1})$ is an $\varepsilon$-FSE in $G(\hat{\rho})$ and $U_{P}(\mathcal{X}(\hat{\rho}),\mathbf{1};f_{\hat{\rho}})\geq\text{OPT}-\varepsilon$.
@@
Let $\tilde{\rho}=\frac{\lfloor\rho^{\star}\cdot\lfloor 1/\delta\rfloor\rfloor}{\lfloor 1/\delta\rfloor}$ be the closest grid point less than or equal to $\rho^{\star}$ that is evaluated by the algorithm in Line 3. Note that $|\tilde{\rho}-\rho^{\star}|\leq\delta$ due to the grid resolution. Additionally, let $\tilde{\mathbf{x}}$ be some arbitrary quality profile that could have been found in Line 5 when evaluating $\tilde{\rho}$. Namely, $\left\|\tilde{\mathbf{x}}-\mathcal{X}(\tilde{\rho})\right\|_{2}\leq\delta$.
@@
We start from demonstrating that $\tilde{\mathbf{x}}$ will pass the $\eta$-FSE check in Line 6 (recall that the algorithm is executed with $\eta=\frac{\varepsilon}{4}$). Recall that we demonstrated in Proposition 16 that $\mathcal{X}(\cdot)$ is Lipschitz continuous with a Lipschitz constant $L$. By the Lipschitz property and the triangle inequality, we have:
@@
Since $(\mathcal{X}(\rho^{\star}),\mathbf{1})$ is an FSE in $G(\rho^{\star})$, it is also trivially 00-FSE. We now want to use Proposition 20 to show that $\tilde{\mathbf{x}}$ is approximate FSE in $G(\tilde{\rho})$ within the required tolerance. To do so, we need the inequality $0.5Y_{LB}>\sqrt{n}\delta=\frac{\sqrt{n}\varepsilon}{A}$ to hold. Remember that if $\varepsilon>\mu X_{UB}$, then the algorithm will return an $\varepsilon$-FSE regardless of our analysis, so we can continue by assuming that $\varepsilon\leq\mu X_{UB}$. As we constructed $A$ such that $A\geq\frac{\sqrt{n}\mu X_{UB}}{0.5Y_{LB}}$, then we have $0.5Y_{LB}\geq\frac{\sqrt{n}\mu X_{UB}}{A}\geq\frac{\sqrt{n}\varepsilon}{A}$, which implies that the condition holds. Thus, employing Proposition 20, we can conclude that $\tilde{\mathbf{x}}$ is an $M_{1}(1+L)\delta$-FSE in $G(\tilde{\rho})$. As we have $\delta=\frac{\varepsilon}{A}$ and we set $A\geq 4M_{1}(1+L)$, we have that $\tilde{\mathbf{x}}$ is an $\frac{\varepsilon}{4}$-FSE in $G(\tilde{\rho})$. Therefore, $\tilde{\mathbf{x}}$ satisfies the condition in Line 6, and the algorithm will find some feasible output.
Now, we show that the evaluated utility $U_{P}(\tilde{\mathbf{x}},\mathbf{1};f_{\tilde{\rho}})$ is close to OPT. For that, we once again use the validity of the condition $0.5Y_{LB}>\sqrt{n}\delta$ (from the same reasoning as above), and employ Proposition 22 to get:
$U_{P}(\tilde{\mathbf{x}},\mathbf{1};f_{\tilde{\rho}})\geq U_{P}(\mathcal{X}(\rho^{\star}),\mathbf{1};f_{\rho^{\star}})-M_{2}(1+L)\delta\geq\text{OPT}-\frac{1}{2}\varepsilon-M_{2}(1+L)\delta.$
As we constructed $A$ such that $A\geq 4M_{2}(1+L)$, we have that
$U_{P}(\tilde{\mathbf{x}},\mathbf{1};f_{\tilde{\rho}})\geq\text{OPT}-\frac{1}{2}\varepsilon-M_{2}(1+L)\frac{\varepsilon}{A}\geq\text{OPT}-\frac{1}{2}\varepsilon-\frac{1}{4}\varepsilon=\text{OPT}-\frac{3}{4}\varepsilon.$
@@
To wrap up the approximation guarantees, we need to show that the returned induced profile, denoted as $\mathcal{X}(\hat{\rho})$, is $\varepsilon$-FSE and that $U_{P}(\mathcal{X}(\hat{\rho}),\mathbf{1};f_{\hat{\rho}})\geq\text{OPT}-\varepsilon$. Notice that this is not immediate from the above analysis, as the corresponding evaluated profile, denoted as $\hat{\mathbf{x}}$, is not necessary the induced profile from the output $\hat{\rho}$. To overcome this mismatch, we once again use Proposition 20 and Proposition 22.
Starting from the $\varepsilon$-FSE guarantee, remember that $\left\|\hat{\mathbf{x}}-\mathcal{X}(\hat{\rho})\right\|_{2}\leq\delta$. Thus, Proposition 20 implies that given that $\hat{\mathbf{x}}$ is an $\eta$-FSE with respect to $\hat{\rho}$, we have that $\mathcal{X}(\hat{\rho})$ is an $\eta+M_{1}\delta$-FSE with respect to $\hat{\rho}$. As $M_{1}\delta=M_{1}\frac{\varepsilon}{A}\leq\frac{1}{4}\varepsilon$, we have that $\mathcal{X}(\hat{\rho})$ is an $\frac{1}{2}\varepsilon$-FSE with respect to $\hat{\rho}$. Moving on to the utility guarantee, give that $U_{P}(\hat{\mathbf{x}},\mathbf{1};f_{\hat{\rho}})\geq\text{OPT}-\frac{3}{4}\varepsilon$, Proposition 22 implies that $U_{P}(\mathcal{X}(\hat{\rho}),\mathbf{1};f_{\hat{\rho}})\geq U_{P}(\hat{\mathbf{x}},\mathbf{1};f_{\hat{\rho}})-M_{2}\delta$. As we set $A$ such that $A\geq 4M_{2}(1+L)$, we have that $M_{2}\frac{\varepsilon}{A}\leq\frac{1}{4}\varepsilon$, and thus $U_{P}(\mathcal{X}(\hat{\rho}),\mathbf{1};f_{\hat{\rho}})\geq\text{OPT}-\varepsilon$.

### v2

@@ C.3 Lipschitz Continuity Results Used in the Proof of Algorithm˜1
Let $\rho,\rho^{\prime}\in[0,1]$ such that $\left|\rho-\rho^{\prime}\right|\leq\delta$, and $\mathbf{x},\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$ and one of those profiles is an ESE in either $G(\rho)$ or $G(\rho^{\prime})$. Assuming $0.5Y_{LB}>\sqrt{n}\delta$, if $\mathbf{x}$ is an $\eta$-FSE in $G(\rho)$, then $\mathbf{x}^{\prime}$ is a $(\eta+B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)\delta)$-FSE in $G(\rho^{\prime})$.
Proof of Lemma˜24.
Define $\bar{\mathbf{x}}\in\{\mathbf{x},\mathbf{x}^{\prime}\}$ to be the ESE profile among the two. Let $\phi_{k}(\mathbf{z},\sigma)$ denote the maximum utility gain player $k$ can achieve by unilaterally deviating from the profile $(\mathbf{z},\mathbf{1})$ when the platform’s revenue sharing parameter is $\sigma$. To apply Lemma 22, we need to establish bounds on other players’ contributions in the profiles $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
From Lemma 18, we know that $\left\|\bar{\mathbf{x}}_{-k}\right\|\geq\sum_{i}x_{i}^{\min}-\max_{k}x_{k}^{\min}\eqqcolon Y_{LB}$ for any player $k$ and any $\rho\in[0,1]$. Additionally, we know from Lemma 13 that $\left\|\bar{\mathbf{x}}_{-k}\right\|\leq\sum_{j=1}^{n}x_{j}^{\max}-\min_{k}x_{k}^{\max}\eqqcolon Y_{UB}$.
We now derive lower bounds for $\left\|\mathbf{x}_{-k}\right\|,\left\|\mathbf{x}^{\prime}_{-k}\right\|$ as well via the assumptions that $\left\|\mathbf{x}-\bar{\mathbf{x}}\right\|_{2},\left\|\mathbf{x}^{\prime}-\bar{\mathbf{x}}\right\|_{2}\leq\delta$ and $0.5Y_{LB}>\sqrt{n}\delta$. Specifically, for any player $k$:
$\left|\left\|\bar{\mathbf{x}}_{-k}\right\|-\left\|\mathbf{x}_{-k}\right\|\right|\leq\left\|\bar{\mathbf{x}}_{-k}-\mathbf{x}_{-k}\right\|\leq\sqrt{n-1}\left\|\bar{\mathbf{x}}_{-k}-\mathbf{x}_{-k}\right\|_{2}\leq\sqrt{n}\delta,$
and the same holds for $\mathbf{x}^{\prime}$. Since $\left\|\bar{\mathbf{x}}_{-k}\right\|\geq Y_{LB}$ and $0.5Y_{LB}\geq\sqrt{n}\delta$, we have $\left\|\mathbf{x}_{-k}\right\|\geq Y_{LB}-\sqrt{n}\delta>0.5Y_{LB}\eqqcolon\tilde{Y}_{LB}>0$. Additionally, we have $\left\|\mathbf{x}_{-k}\right\|\leq Y_{UB}+\sqrt{n}\delta<2Y_{UB}\eqqcolon\tilde{Y}_{UB}$. These bounds hold for $\mathbf{x}^{\prime}$ as well. Notice that the validity of these bounds does not depend on which of the two profiles is the ESE.
Now, we analyze the deviation gain of any player from the profile $\mathbf{x}^{\prime}$. By the assumption that $\mathbf{x}$ is $\eta$-FSE in $G(\rho)$, we know that $\phi_{k}(\mathbf{x},\rho)\leq\eta$ for all players $k$. Applying the triangle inequality, we have:
$\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\leq\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|+\phi_{k}(\mathbf{x},\rho).$
Now, applying Lemma 22 with $\tilde{Y}_{LB},\tilde{Y}_{UB}$ as the bounds for other players’ contributions, we get:
@@
$\left|\phi_{k}(\mathbf{x},\rho)-\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\right|\leq\delta\cdot\left(\tilde{B}(1,1,0,1,2-\gamma)+\tilde{B}(0,(2+\alpha),0.5,2,3-\gamma)+\tilde{B}(0,2,0,1,1-\gamma)\right),$
where $\tilde{B}$ is defined similarly to $B$ but with $\tilde{Y}_{LB},\tilde{Y}_{UB}$ instead of $Y_{LB},Y_{UB}$. Thus, we have:
$\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\leq\delta\cdot\left(\tilde{B}(1,1,0,1,2-\gamma)+\tilde{B}(0,(2+\alpha),0.5,2,3-\gamma)+\tilde{B}(0,2,0,1,1-\gamma)\right)+\eta.$
We now want to simplify the bracketed term. First, notice that
$\tilde{B}(1,1,0,1,2-\gamma)+\tilde{B}(0,(2+\alpha),0.5,2,3-\gamma)+\tilde{B}(0,2,0,1,1-\gamma)$
$\leq\tilde{B}(1,1,0,2,3-\gamma)+\tilde{B}(0,(2+\alpha),0.5,2,3-\gamma)+\tilde{B}(0,2,0,3,3-\gamma)$
$\leq\tilde{B}(1,(1+2+\alpha+2),0.5,3,3-\gamma)$
$=\tilde{B}(1,(5+\alpha),0.5,3,3-\gamma)$
$\leq B(1,(5+\alpha)\cdot 2^{3}\cdot 2^{3-\gamma},0.5,3,3-\gamma)$
$=B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma),$
where the second to last inequality follows from the definition of $B$ and the fact that $\tilde{Y}_{LB}=0.5Y_{LB}$ and $\tilde{Y}_{UB}=2Y_{UB}$. After this simplification, we get that:
$\phi_{k}(\mathbf{x}^{\prime},\rho^{\prime})\leq B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)\cdot\delta+\eta,$
which means that $\mathbf{x}^{\prime}$ is a $(\eta+B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)\delta)$-FSE in $G(\rho^{\prime})$, completing the proof of Lemma˜24. ∎
Lemma 25.
Let $\rho,\rho^{\prime}\in[0,1]$ such that $\left|\rho-\rho^{\prime}\right|\leq\delta$, and $\mathbf{x},\mathbf{x}^{\prime}$ such that $\left\|\mathbf{x}-\mathbf{x}^{\prime}\right\|_{2}\leq\delta$ and one of those profiles is an ESE in either $G(\rho)$ or $G(\rho^{\prime})$. Assuming $0.5X_{LB}\geq\sqrt{n}\delta$, it holds that
$\left|U_{P}(\mathbf{x},\mathbf{1};f_{\rho})-U_{P}(\mathbf{x}^{\prime},\mathbf{1};f_{\rho^{\prime}})\right|\leq B(0,4(1+\gamma),0.5,1,1-\gamma)\delta.$
Proof of Lemma˜25.
Without loss of generality, we assume $\mathbf{x}$ is the unique ESE in $G(\rho)$. Following the same steps as in the proof of Lemma˜24, we get that $0.5X_{LB}\leq\left\|\mathbf{x}\right\|\leq 2X_{UB}$, as well as $\left\|\mathbf{x}^{\prime}\right\|$ bounds. Next, we apply Lemma 23 with those bounds to get:
$\left|U_{P}(\mathbf{x},\mathbf{1};f_{\rho})-U_{P}(\mathbf{x}^{\prime},\mathbf{1};f_{\rho^{\prime}})\right|$ $\leq\tilde{B}(0,1+\gamma,0.5,1,1-\gamma)\cdot\delta$
$\leq B(0,4(1+\gamma),0.5,1,1-\gamma)\cdot\delta,$
which completes the proof of Lemma˜25. ∎
Lemma 26.
Let $f(x_{i})=\frac{\mu(x_{i}+\left\|\mathbf{x}_{-i}\right\|)^{\gamma}x_{i}}{x_{i}+(1+\alpha)\left\|\mathbf{x}_{-i}\right\|}-c_{i}(x_{i})$ be the utility function for player $i$ under no sharing. Assume we have a profile $\mathbf{x}$ such that $\left\|\mathbf{x}-\mathcal{X}(\rho)\right\|_{2}\leq\delta$, where $\mathcal{X}(\rho)$ is the ESE in $G(\rho)$. Assuming that $0.5Y_{LB}>\delta\sqrt{n}$, $f(x_{i})$ is Lipschitz continuous in $x_{i}$ with Lipschitz constant $B(1,5\cdot 2^{5-\gamma},0,2,3-\gamma)$.
Proof of Lemma˜26.
@@
Following the logic in the proof of Lemma˜24, we have that $\left\|\mathbf{x}_{-i}\right\|$ is bounded as $0.5Y_{LB}<\left\|\mathbf{x}_{-i}\right\|<2Y_{UB}$. Let $y=\left\|\mathbf{x}_{-i}\right\|$ for simplicity. We compute the derivative of $f(x_{i})$ using the quotient rule:
@@
Using our established bounds $y\geq 0.5Y_{LB}$, $y\leq 2Y_{UB}$, and $x_{i}\leq x_{i}^{\max}$:
$\left|f^{\prime}(x_{i})\right|$ $\leq\mu\Bigg[\frac{\gamma(0.5Y_{LB})^{\gamma-1}(x_{i}^{\max})^{2}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{\gamma(0.5Y_{LB})^{\gamma-1}x_{i}^{\max}(1+\alpha)(2Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(x_{i}^{\max}+2Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}+\frac{(x_{i}^{\max}+2Y_{UB})^{\gamma}(1+\alpha)(2Y_{UB})}{[(1+\alpha)(0.5Y_{LB})]^{2}}$
$+\frac{(2Y_{UB})^{\gamma}x_{i}^{\max}}{[(1+\alpha)(0.5Y_{LB})]^{2}}\Bigg]+\left|c_{i}^{\prime}(x_{i}^{\max})\right|$
$\leq\tilde{B}(0,1,0,2,3-\gamma)+\tilde{B}(0,1,0,2,3-\gamma)+\tilde{B}(0,1,0,\gamma+1,2)+\tilde{B}(0,1,0,\gamma+1,2)$
$+\tilde{B}(0,1,0,\gamma+1,2)+c_{i}^{\prime}(x_{i}^{\max})$
$\leq\tilde{B}(1,5,0,2,3-\gamma)$
$\leq B(1,5\cdot 2^{5-\gamma},0,2,3-\gamma).$
This completes the proof of Lemma˜26. ∎
C.4 Strong Monotonicity and Lipschitz Continuity of ESE
Lemma 27.
Proof of Lemma˜27.
We’ll extend the monotonicity result from Lemma˜3 to show that with $\beta$-strongly convex cost functions, the game becomes strongly monotone.
Recall from the proof of Lemma˜3 that:
We already established that the first term is non-positive:
Summing over all players:
Therefore:
Lemma 28.
Assume that all the cost functions $c_{i}(x_{i})$ are $\beta$-strongly convex. Then, $\mathcal{X}(\cdot)$ is Lipschitz continuous with respect to the $\ell_{2}$-norm on the interval $[0,1]$, with Lipschitz constant at most $B(0,1/\beta,0.5,1,2-\gamma)$.
Proof of Lemma˜28.
Step 1: Characterizing the equilibrium
Denote $K(\rho)=\frac{\mu(1+\alpha\rho)}{1+\alpha}$. The utility for creator $i$ under enforced full sharing ($\mathbf{s}=\mathbf{1}$) is given by Lemma˜12:
Let us define $g_{i}(\mathbf{x},\rho)=\frac{\partial U_{i}}{\partial x_{i}}=K(\rho)\left\|\mathbf{x}\right\|^{\gamma-2}\left[\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|\right]-c_{i}^{\prime}(x_{i})$. Recall that the FOC for creator $i$ in the ESE is $g_{i}(\mathcal{X}(\rho),\rho)=0$.
Step 2: Continuous differentiability of $\mathcal{X}(\cdot)$
We first verify that $\mathbf{g}(\mathbf{x},\rho)$ is continuously differentiable in both $\mathbf{x}$ and $\rho$:
  • $K(\rho)=\frac{\mu(1+\alpha\rho)}{1+\alpha}$ is clearly $C^{1}$ in $\rho$.
  • From Lemma 15 and the arguments in the proof of Theorem 5, we know that $\mathcal{X}_{i}(\rho)>0$ for all $i\in[n]$. Consequently, terms like $\left\|\mathbf{x}\right\|^{\gamma-2}$ are well-defined and $C^{1}$ in a neighborhood of $\mathcal{X}(\rho)$ and are continuously differentiable as a composition of $C^{1}$ functions.
This condition of strong monotonicity for a continuously differentiable map $\mathbf{g}$ implies that the symmetric part of its Jacobian, $J_{S}(\mathbf{x},\rho)=\frac{1}{2}(J_{x}(\mathbf{x},\rho)+J_{x}(\mathbf{x},\rho)^{T})$, is negative definite. This, in turn, implies that the Jacobian $J_{x}(\mathcal{X}(\rho),\rho)$ itself is invertible. To see why, assume for contradiction that $J_{x}(\mathcal{X}(\rho),\rho)$ is not invertible. Then, there exists a non-zero vector $\mathbf{u}$ such that $J_{x}(\mathcal{X}(\rho),\rho)\mathbf{u}=\mathbf{0}$. This would lead to:
Step 3: Bounding the derivative norm
First, we bound $\left\|J_{x}(\mathcal{X}(\rho),\rho)^{-1}\right\|_{2}$. Using standard results from variational inequalities (e.g., [38]), we know that strong monotonicity of $\mathbf{g}(\mathbf{x},\rho)$ implies that the symmetric part of the Jacobian $J_{S}(\mathcal{X}(\rho),\rho)$ is negative definite with the following property: for any non-zero vector $\mathbf{v}$,
Lemma˜18 ensures that $X_{LB}\leq\left\|\mathcal{X}(\rho)\right\|$. Additionally, Lemma˜13 suggests that $\left\|\mathcal{X}(\rho)\right\|\leq X_{UB}$. Using the fact that $\gamma-2<0$, we have $\left\|\mathcal{X}(\rho)\right\|^{\gamma-2}\leq(X_{LB})^{\gamma-2}$. Also, $\gamma\mathcal{X}_{i}(\rho)+\left\|\mathcal{X}_{-i}(\rho)\right\|\leq X_{UB}$ and $\frac{\mu\alpha}{1+\alpha}\leq\mu$. Therefore:
Using the Euclidean norm:
Combining our bounds:
Since $\mathcal{X}(\cdot)$ is continuously differentiable on the compact interval $[0,1]$, its derivative is bounded. By the Mean Value Theorem, for any $\rho,\rho^{\prime}\in[0,1]$:
$\left\|\mathbf{x}^{\star}(\rho)-\mathbf{x}^{\star}(\rho^{\prime})\right\|_{2}\leq L\cdot\left|\rho-\rho^{\prime}\right|,$
where $L=\sup_{\rho\in[0,1]}\left\|\frac{d\mathbf{x}^{\star}}{d\rho}(\rho)\right\|_{2}\leq\frac{\mu\sqrt{n}}{\beta}\frac{X_{UB}}{(X_{LB})^{2-\gamma}}\leq B(0,1/\beta,0.5,1,2-\gamma)$. The proof of Lemma˜28 is now complete. ∎
C.5 Proofs of Section 4’s Main Results
See 8
Proof of Theorem˜8.
@@ C.5 Proofs of Section 4’s Main Results
Before starting the proof, we address a special case of the choice of a large $\varepsilon$. Although we don’t expect the algorithm to be executed with such a large $\varepsilon$, we want to show that the algorithm will prevail in this case as well. Notice that if $\varepsilon>\mu(1+X_{UB})$, then no matter which output the algorithm returns, it will satisfy the approximation guarantees:
@@
  • First, notice that the utilities of the players are bounded by $\mu(1+X_{UB})$. Thus, no player can have a deviation that is larger than $\mu(1+X_{UB})$, and the algorithm must return an $\varepsilon$-FSE.
@@
  • In the same line of reasoning, the platform revenue is also bounded by $\mu(1+X_{UB})$, thus, the algorithm will return a solution with revenue not smaller than $\text{OPT}-\varepsilon$.
@@
Given these reasons, we limit our attention to the case when $\varepsilon\leq\mu X_{UB}$.
Part 1: Approximation guarantees
@@
First, let us specify the choice of the constant $\delta$ in the algorithm. We set $\delta=\frac{\varepsilon}{A}$, where $A=4B(1,(5+\alpha)\cdot 2^{6-\gamma},0.5,3,3-\gamma)\cdot(1+B(0,1/\beta,0.5,1,2-\gamma))$. The reasons for this choice will become clear in the next paragraphs.
[…]
@@
Let $\tilde{\rho}=\frac{\lfloor\rho^{\star}\cdot\lfloor 1/\delta\rfloor\rfloor}{\lfloor 1/\delta\rfloor}$ be the closest grid point less than or equal to $\rho^{\star}$ that is evaluated by the algorithm in Line 4. Note that $|\tilde{\rho}-\rho^{\star}|\leq\delta$ due to the grid resolution. Additionally, let $\tilde{\mathbf{x}}$ be some arbitrary quality profile that could have been found in Line 6 when evaluating $\tilde{\rho}$. Namely, $\left\|\tilde{\mathbf{x}}-\mathcal{X}(\tilde{\rho})\right\|_{2}\leq\delta$.
@@
We start from demonstrating that $\tilde{\mathbf{x}}$ will pass the $\eta$-FSE check in Line 7 (recall that the algorithm is executed with $\eta=\frac{\varepsilon}{4}$). Recall that we demonstrated in Lemma˜28 that $\mathcal{X}(\cdot)$ is Lipschitz continuous with a Lipschitz constant $L=B(0,1/\beta,0.5,1,2-\gamma)$. By the Lipschitz property and the triangle inequality, we have:
@@
[…]
$\geq\text{OPT}-\frac{1}{2}\varepsilon-B(0,4(1+\gamma),0.5,1,1-\gamma)(1+L)\delta.$
As we constructed $A$ such that $A\geq 4B(0,4(1+\gamma),0.5,1,1-\gamma)(1+L)$, we have that
$U_{P}(\tilde{\mathbf{x}},\mathbf{1};f_{\tilde{\rho}})\geq\text{OPT}-\frac{1}{2}\varepsilon-B(0,4(1+\gamma),0.5,1,1-\gamma)(1+L)\frac{\varepsilon}{A}\geq\text{OPT}-\frac{1}{2}\varepsilon-\frac{1}{4}\varepsilon=\text{OPT}-\frac{3}{4}\varepsilon.$
@@
[…]